# A dark matter hail: Detecting macroscopic dark matter with asteroids, planetary rings, and craters

### v1

@@ IV Orbit evaporation
$k\geq\left(\frac{v_{o}}{v_{DM}}\right)^{2}\left(\frac{M}{m}\right)^{2}\left(\delta-1\right)^{2}~{}.$ (22)
@@
$m\geq\frac{16\pi}{9}\left(\delta-1\right)^{2}\frac{v_{o}^{2}}{v_{DM}^{3}}\frac{R^{4}}{\tau}\frac{\rho_{o}^{2}}{\rho_{DM}}~{}.$ (23)
If we have $k>1$ and Eqs. 18 and 23 are satisfied for $\delta=\sqrt{2}$, we can consider the dark matter scenario constrained. For consistency, we should also insist that the object is not trivially destroyed in this process, using the calculations of the previous section. We will assume that if one collision does not destroy the object, then many subsequent ones will not either—this may not be a safe assumption, but if the object is destroyed then we can anyway still consider the scenario constrained.
@@
As it turns out, the most interesting large objects to consider are Kuiper belt objects. We did not discuss Kuiper belt objects in the previous section since they offered no real advantage over the S-type asteroid destruction constraints. Here however the key advantage is the comparatively low orbital speed, only about $5$ km/s. This allows even larger Kuiper belt objects to be evaporated within the lifetime of the solar system. Kuiper belt objects have been observed with radii spanning approximately $1-1000$ km, the largest of which is Pluto [49]. Importantly, Kuiper belt objects above $40$ km or so in radius are thought to be pristine remnants of the formation of the solar system [49].
These constraints could be improved again by considering Oort cloud objects [79, 80], which would have have yet smaller orbital speeds than Kuiper belt objects. However, since observations of the Oort cloud is even today limited to indirect observations of comet trajectories, these limits will not be included here.
Saturn’s rings again turn out to be interesting, since they have a range of orbital velocities spanning roughly $2-25$ km/s [66]. The strongest constraints thus come from the lower part of this range, in the outer rings. One issue here is that as previously discussed, the particles in Saturn’s rings are constantly colliding with each other and thus distributing the extra energy they would get from the dark matter collisions. However, since all the particles should be ‘heating up’ roughly at the same rate, the effect of many self-collisions should ultimately not stop the total mass of the rings from evaporating. Again, a more precise constraint would require long timespan numerical simulations of the evolution of Saturn’s rings, including the incident dark matter population—and ideally, including the effect of the dark matter wind, which should only improve these bounds.
Ultimately the constraints for both the Kuiper belt objects and the evaporation of Saturn’s rings are plotted in Fig. 1. For clarity, the constraints for the evaporation of the asteroid belt, other ring systems, or other large bodies are not plotted.
@@ V Earth cratering
Now we investigate constraints on macroscopic dark matter from the frequency of collisions with Earth and the formation of craters. The rate of impacts on Earth for a given dark matter mass have also been indicated on Fig. 1 with ticks along the bottom axis. While one might be concerned that these rates seem quite high, it ultimately is not easy to constrain the dark matter using Earth impacts. These constraints should be expected to overlap significantly with the asteroid constraints, since one of the earlier methods for determining if an asteroid was destroyed was by considering the size of craters—if the crater would be larger than the asteroid, it was considered destroyed [35, 34].
@@
where $\ell$ is the radius of the crater. Of course there is a wide variance in cratering from factors such as the ground material, density of the meteor, and the angle and velocity of impact, so this estimate should be taken as a useful guide for order-of-magnitude estimations.
@@
$r\geq 0.89~{}\mathrm{km}~{}\left(\frac{\ell}{km}\right)^{0.80}\left(\frac{\rho_{Earth}}{3\mathrm{gcm}^{-3}}\right)^{-1/2}\left(\frac{v_{DM}}{220\mathrm{km/s}}\right)^{-1}~{}.$ (25)
To estimate the constraints, however, we need empirical data about the rate of asteroid collisions with Earth. In general, this is extremely complicated, because the active geological history of the Earth makes detailed evidence scant. While we can learn a significant amount from the hundreds-or-so of known craters [42], estimating the overall rate of impacts from these surviving craters is more difficult, since one would need to model the number of events which didn’t lead to craters discoverable today. Such modeling is beyond the scope of this work, and possibly an impossible challenge given the rich geological history of Earth. This makes it very difficult to place confident constraints on macroscopic dark matter purely from the numbers of craters on active geological bodies like Earth.
@@
One interesting way to improve the constraints could be to carefully account for the dark matter wind. The Solar System is at the moment moving in the direction of the Cygnus constellation, at around $40^{\circ}$ declination. This means that the North pole of the Earth, as well as many other bodies whose spin axes are approximately perpendicular to the plane of the Solar System, would see a larger flux of dark matter impacts at present day than their South poles. If there is no difference in the present cratering rate across the North and South poles of objects like the moon, stronger constraints on the dark matter could perhaps be placed. This is somewhat stymied by the fact that the dark matter wind will be averaged out over $250$Myr timescales, so these constraints would need to focus on timescales closer to the order of $10$Myr.
@@ VI Results and conclusions
If the dark matter is composed of macroscopic objects, it could manifest as a ‘hail’ impacting on objects in our Solar System. This work presents order-of-magnitude estimates for using a number of Solar System bodies—asteroids, planetary rings, and terrestrial cratering—in constraining macroscopic dark matter. Our new dark matter constraints are plotted along with existing constraints in Fig. 1. Previous constraints come from a wide variety of sources, spanning cosmology, astrophysics, geology, and more traditional dark matter direct detection. The constraints included on this plot (some of them hidden behind new constraints) include the cosmic microwave background (CMB) [87, 88], the Milky Way satellites (MW) [89], microlensing [90, 91], white dwarves [92], neutron stars [91], human death [93], fireballs [94], meteors, radar [95], ancient mica [96], Ohya [97], Skylab [97], Chicago [98], DEAP [99], XQC [100], DAMA [101, 100], CRESST [100], CDMS [100], XENON1T [102], and LZ [103]. These constraints, along with the plotting code, were compiled and published by Ciaran O’Hare at Ref. [32].

### v2

@@ IV Orbit evaporation
$\sqrt{k}\geq\frac{1}{0.89}\frac{v_{o}}{v_{\rm rel}}\frac{M}{m}(\delta-1)~.$ (21)
@@
$m\geq\frac{16\pi}{9(0.89^{2})}\left(\delta-1\right)^{2}\frac{v_{o}^{2}}{v_{\rm rel}^{3}}\frac{R^{4}}{\tau}\frac{\rho_{o}^{2}}{\rho_{DM}}~.$ (22)
If we have $k>1$ and Eqs. 17 and 22 are satisfied for $\delta=\sqrt{2}$, we can consider the dark matter scenario constrained. For consistency, we should also insist that the object is not trivially destroyed in this process, using the calculations of the previous section. We will assume that if one collision does not destroy the object, then many subsequent ones will not either—this may not be a safe assumption, but if the object was to be destroyed by multiple collisions, then we could anyways still consider the scenario constrained.
@@
As it turns out, the most interesting large objects to consider are Kuiper belt objects. We did not discuss Kuiper belt objects in the previous section since they offered no real advantage over the S-type asteroid destruction constraints. Here however the key advantage is the comparatively low orbital speed of around $5$ km/s. This allows even larger Kuiper belt objects to be evaporated within the lifetime of the Solar System. Kuiper belt objects have been observed with radii spanning approximately 110001-1000 km, the largest of which is Pluto [49]. Importantly, Kuiper belt objects above $40$ km or so in radius are thought to be pristine remnants of the formation of the Solar System. For the lifetimes of Kuiper belt objects, we follow Ref. [49], who infer the collisional lifetimes of differently-sized Kuiper belt objects from the location of the kink in their mass spectrum.
These constraints could be improved again by considering Oort cloud objects [80, 81], which would have have yet smaller orbital speeds than Kuiper belt objects. However, since observations of the Oort cloud are even today limited to indirect observations of comet trajectories, these limits will not be included here.
Saturn’s rings again turn out to be interesting, since they have a range of orbital velocities spanning roughly 2252-25 km/s [66]. The strongest constraints thus come from the lower part of this range, in the outer rings. One issue here is that as previously discussed, the particles in Saturn’s rings are constantly colliding with each other and thus distributing the extra energy they would get from the dark matter collisions. However, since all the particles should be ‘heating up’ roughly at the same rate, the effect of many self-collisions should ultimately not stop the total mass of the rings from evaporating. Again, a more precise constraint would require long timespan numerical simulations of the evolution of Saturn’s rings, including the incident dark matter population.
The constraints for both the Kuiper belt objects and the evaporation of Saturn’s rings are plotted in Fig. 1. For clarity, the constraints for the evaporation of the asteroid belt, other ring systems, or other large bodies are not shown, since they are superseded by the plotted constraints.
@@ V Earth cratering
Now we investigate constraints on macroscopic dark matter from the frequency of collisions with Earth and the formation of craters. The rate of impacts on Earth for a given dark matter mass have also been indicated on Fig. 1 with ticks along the bottom axis. While one might be concerned that these rates seem quite high, it ultimately is not easy to constrain the dark matter using Earth impacts. These constraints should be expected to overlap significantly with the asteroid constraints, since one of the first methods historically for determining if an asteroid was destroyed was to consider the size of craters—if the crater would be larger than the asteroid, it was considered destroyed [35, 34].
@@
where $\ell$ is the radius of the crater. Of course there is a wide variance in cratering when considering factors like the ground material, density of the meteor, and the angle and velocity of impact, so this estimate should be taken only as a useful guide for order-of-magnitude estimations.
@@
$r\geq 0.89~\mathrm{km}~\left(\frac{\ell}{km}\right)^{0.80}\left(\frac{\rho_{\oplus}}{3\mathrm{gcm}^{-3}}\right)^{-1/2}\left(\frac{v_{DM}}{220\mathrm{km/s}}\right)^{-1}~,$ (24)
where $\rho_{\oplus}$ is the density of the Earth’s surface material. To estimate the constraints, however, we need empirical data about the rate of asteroid collisions with Earth. In general, this is extremely complicated, because the active geological history of the Earth makes detailed evidence scant. While we can learn a significant amount from the hundreds-or-so of known craters [42], estimating the overall rate of impacts from these surviving craters is more difficult, since one would need to model the number of events which didn’t lead to craters discoverable today. Such modeling is beyond the scope of this work, and possibly an impossible challenge given the rich geological history of Earth. This makes it very difficult to place confident constraints on macroscopic dark matter purely from the numbers of craters on active geological bodies like Earth.
@@
One interesting way to improve the constraints could be to carefully account for the dark matter wind. As discussed earlier, the Solar System is at the moment moving in the direction of the Cygnus constellation, at around $40^{\circ}$ declination. This means that the North pole of the Earth, as well as many other bodies whose spin axes are approximately perpendicular to the plane of the Solar System, would see a larger flux of dark matter impacts at present day than their South poles. If there is no difference in the present cratering rate across the North and South poles of objects like the moon, stronger constraints on the dark matter could perhaps be placed. This is somewhat stymied by the fact that the dark matter wind will be averaged out over $230$ Myr timescales, so these constraints would need to focus on timescales closer to the order of $10$ Myr.
@@ VI Results and conclusions
If the dark matter is composed of macroscopic objects, it could manifest as a ‘hail’ impacting on objects in our Solar System. This work presents order-of-magnitude estimates for using a number of Solar System bodies—asteroids, planetary rings, and terrestrial cratering—in constraining macroscopic dark matter. Our new dark matter constraints are plotted along with existing constraints in Fig. 1. Previous constraints come from a wide variety of sources, spanning cosmology, astrophysics, geology, and more traditional dark matter direct detection. The constraints included on this plot (some of them hidden behind new constraints) include the cosmic microwave background (CMB) [88, 89], Milky Way satellites (MW) [90], microlensing [91, 92], white dwarves [93], neutron stars [92], human death [94], fireballs [95], meteor radar data [96], ancient mica [97], Ohya [98], Skylab [98], Chicago [99], DEAP [100], XQC [101], DAMA [102, 101], CRESST [101], CDMS [101], XENON1T [103], and LZ [104]. These constraints, along with the plotting code, were compiled and published by Ciaran O’Hare in Ref. [32].